OpenStudy (anonymous):

get a medal What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0 {(31/8, 25/8)} {(25/8, 31/8)} {(45/2, 31/2)} {(31/2, 45/2)}

OpenStudy (anonymous):

this first 1 i think

OpenStudy (anonymous):

hhhhmmmmmmmm i need u to be sure

OpenStudy (anonymous):

2) x + y - 7 = 0>>>>>>>x=-y+7 .........._(3) substitute (3) in (1): 5(-y+7)-3y=10 -5y+35-3y=10 -8y=-25 y=25/8 >>>> (4) substitute (4) in (3): x=(-25/8)+7=31/8 so,the answer is :{(31/8, 25/8)